# Reasoning Models Are More Easily Gaslighted Than You Think

### v1

@@ 3.2. Reasoning Gaslighting Benchmark Construction
To systematically evaluate reasoning model robustness against adversarial user interventions, we introduce GaslightingBench-R (“R” indicates Reasoning), a new benchmark designed to assess the susceptibility of the reasoning models to gaslighting-style negation prompts. GaslightingBench-R is constructed from the above three multimodal reasoning datasets: MMMU, MathVista, and CharXiv. These datasets are selected for their domain diversity, multimodal grounding, and established utility in evaluating advanced reasoning models. Our goal in constructing GaslightingBench-R is twofold: (1) Ensure diversity across reasoning categories to support broad evaluation coverage. (2) Prioritize samples where models exhibit a shift from correct to incorrect answers after gaslighting prompts, ensuring high signal for susceptibility. We adopt a model-driven selection pipeline informed by the gaslighting vulnerability of individual samples. Specifically, for each sample in the candidate datasets, we calculate a vulnerability score defined as:
(1) $\text{Score}=\sum_{i=1}^{3}\mathbbm{1}^{(i,\text{before})}-\sum_{i=1}^{3}\mathbbm{1}^{(i,\text{after})},$
@@
where $\mathbbm{1}^{(i,\text{before})}$ and $\mathbbm{1}^{(i,\text{after}}$ indicate whether model $i$ (among the three evaluated: OpenAI o4-mini, Claude-3.7-Sonnet, Gemini-2.5-Flash) produce the correct answer before and after the gaslighting negation, respectively. Samples with higher scores indicate greater susceptibility to gaslighting and are prioritized during sampling. To ensure domain diversity, we stratify selection across categories defined by each dataset. As shown in Figure 2, for MMMU and CharXiv, we uniformly sample from non-overlapping subject categories. MathVista, however, presents a special case: its examples often span multiple overlapping skills (e.g., a single question tagged as both geometry and numeric commonsense). As a result, category counts in MathVista are not strictly balanced, though we aim for representative coverage across reasoning types. GaslightingBench-R is concluded with 21 categories and 1,025 samples after careful curation to reflect both reasoning diversity and gaslighting vulnerability.
@@
MMMU (Yue et al., 2024) MathVista (Lu et al., 2024) CharXiv (Wang et al., 2024)
@@
Figure 3. Performance comparison of reasoning models on GaslightingBench-R before and after gaslighting negation prompts.
4. Result Analysis
Figure 4. Accuracy comparison before and after gaslighting negation across subject categories in the MMMU benchmark. The bars represent model accuracy before (light) and after (dark) gaslighting negation, with the absolute accuracy drop annotated.
Figure 5. Reasoning skill-wise accuracy comparison on the MathVista benchmark before and after gaslighting negation prompts.
Figure 6. Subject-wise accuracy comparison on the CharXiv benchmark before and after gaslighting negation prompts.
4.1. Performance Comparison
@@
As shown in Table 1, we compare the performance of three state-of-the-art reasoning models, OpenAI-o4-mini, Claude-3.7-Sonnet and Gemini-2.5-Flash, before and after the introduction of gaslighting negation prompts across three multimodal reasoning benchmarks. All three models exhibit significant performance degradation after the introduction of gaslighting prompt. The average accuracy drop across the three datasets ranges from 25.6 to 28.8 percent. These results demonstrate that even the most advanced reasoning models struggle to maintain factual consistency and are unable to reliably defend correct answers under adversarial user interventions.
@@
Before gaslighting negation, OpenAI-o4-mini consistently outperforms its counterparts, achieving accuracies of 77.4 on MMMU, 77.1 on MathVista, and 65.2 on CharXiv, resulting in an overall average of 73.2%. Following the gaslighting negation prompt, its performance drops to 52.1, 54.1, and 36.7, respectively, with an average of 47.6%. While this represents a substantial decline of 25.6%, it is the smallest average drop among the three evaluated models. In particular, Gemini-2.5-Flash achieves the highest pre-negation accuracy on MathVista (77.6), but suffers the largest performance decline (35.4) after gaslighting negation, highlighting its high vulnerability and limited robustness in maintaining correct mathematical reasoning under adversarial gaslighting negation.
@@
As shown in Figure 3, we further evaluate all three models on our newly constructed GaslightingBench-R. Compared to their performance on the existing benchmarks, the models exhibit substantially more severe degradation, with accuracy dropping by over 53% on average. In contrast, the average accuracy drops on MMMU, MathVista, and CharXiv range from 25.6% to 28.8%. This discrepancy highlights the unique difficulty of GaslightingBench-R, which is intentionally curated to target model susceptibility. Specifically, samples are selected based on their likelihood to induce belief reversals across multiple models, as described in Section 3.2, thereby maximizing their diagnostic value. Consequently, GaslightingBench-R serves as a focused and stress-tested benchmark for evaluating reasoning model robustness, offering deeper insights into belief collapse and behavioral inconsistency than conventional evaluations.
Figure 7. Qualitative examples of gaslighting-induced answer reversals across reasoning models. The revised answers are accompanied by hallucinated rationales absent from the original explanation, revealing vulnerabilities in the reasoning process.
4.2. Category Analysis and Qualitative Result
@@
Category analysis. Figures 46 present a fine-grained breakdown of accuracy before and after gaslighting prompts for three reasoning models across thematic sub-categories in the MathVista, MMMU, and ChartXiv benchmarks. Figure 4 shows bar plots of model accuracy across six subject categories from the MMMU benchmark: Tech and Engineering, Humanities and Social Sciences, Health and Medicine, Science, Business, and Art and Design. Prior to gaslighting negation, all three models perform relatively well in Business, while exhibiting lower accuracy in Tech and Engineering. Following gaslighting negation prompts, the most substantial performance declines are observed (e.g., –37.3 for OpenAI-o4-mini, –30.0 for Gemini-2.5-flash), indicating that models are particularly fragile in domains requiring precise and high-stakes reasoning. In contrast, the Business category experiences the smallest declines on average (e.g., –16.6 for OpenAI-o4-mini, –18.0 for Claude-3.7-Sonnet), suggesting a relatively stable reasoning trajectory in economically contextual tasks. Figure 5 presents radar plots highlighting the three reasoning models’ performance across six core reasoning skills: geometry, arithmetic, algebraic, logical, statistical reasoning, and numeric commonsense. Before negation, all models exhibit strong performance in geometry and statistical reasoning. After negation, notable performance declines are observed in arithmetic reasoning and numeric commonsense, particularly for OpenAI-o4-mini and Gemini-2.5-Flash. suggesting these areas are particularly vulnerable to gaslighting prompts. These categories may be more susceptible to gaslighting due to their quantitative ambiguity and dependence on intermediate calculations (Geva et al., 2020). Figure 6 provides a radar plot comparing model performance across eight scientific subjects in CharXiv: Mathematics, Physics, Quantitative Biology, Quantitative Finance, Statistics, Computer Science, Economics, and Electrical Engineering and Systems Science. Compared to MMMU and MathVista, while all three models suffer noticeable performance declines after gaslighting across every subject in CharXiv, the degradation is relatively uniform, showing the general vulnerability of gaslighting negation for the real-world scientific chart understanding.
@@
Qualitative result. Figure 7 showcases qualitative examples of gaslighting-induced failures across the three reasoning models, highlighting the models’ behavior before and after exposure to a gaslighting negation prompt. The left figure presents a math-based visual reasoning question, while the right panel involves a conceptual multimodal question about literary interpretation grounded in an image of a castle. In both examples, all the models initially provide the correct answer along with clear, well-reasoned justifications grounded in visual and textual input. However, after receiving a simple negation prompt (e.g., “No, that’s incorrect. Please verify your answer.”), all models revise their answers to incorrect ones. In the math example, all three models abandon their correct initial answer (radius = 5) and switch to an incorrect value (radius = 4) after the negation prompt, despite the visual evidence being clear and objective. Each model introduces flawed reasoning to justify the revision, revealing a lack of belief stability even in straightforward numerical tasks. In the literary example, Claude-3.7-Sonnet and Gemini-2.5-Flash both initially select the correct answer (”Castle Tales”) based on historical and contextual cues, but reverse their decisions to the incorrect ”Retold Tale” after negation. The shift reflects not only factual inconsistency but also hallucinated rationalizations, wherein the model invents supporting arguments for the new answer that were not part of its original reasoning process. This behavior exemplifies the dangers of over-accommodating user feedback at the expense of internal consistency.
5. Conclusion
@@
We have presented the first systematic investigation into the robustness of state-of-the-art reasoning models against adversarial gaslighting negation prompts. Despite their use of chain-of-thought reasoning and test-time scaling, models like OpenAI’s o4-mini, Claude-3.7-Sonnet and Gemini-2.5-Flash exhibit significant belief reversals when challenged with negation prompts after providing correct and well-justified answers. To diagnose this vulnerability more precisely, we introduce GaslightingBench-R, a curated benchmark targeting belief inconsistency in reasoning models. Our results show that GaslightingBench-R elicits even more noticeable failures than standard benchmarks, revealing a critical gap between reasoning transparency and belief stability. These findings call for rethinking evaluation protocols and advancing robustness strategies that go beyond correctness and interpretability to include resilience against adversarial conversational manipulations.

### v2

@@ 3.2 Reasoning Gaslighting Benchmark Construction
To systematically evaluate reasoning model robustness against adversarial user interventions, we introduce GaslightingBench-R (where "R" indicates Reasoning), a new benchmark designed to assess the susceptibility of reasoning models to gaslighting negation attacks. GaslightingBench-R is constructed from the above three multimodal reasoning datasets: MMMU, MathVista, and CharXiv. These datasets are selected for their domain diversity, multimodal grounding, and established utility in evaluating advanced reasoning models. Our goal in constructing GaslightingBench-R is twofold: (1) Ensure diversity across reasoning categories to support broad evaluation coverage. (2) To more clearly demonstrate the gaslighting effect and reduce the evaluation overhead of the subsequent benchmark, we prioritize samples where models exhibit a shift from correct to incorrect answers after gaslighting prompts, ensuring high signal for susceptibility. We adopt a model-driven selection pipeline informed by the gaslighting vulnerability of individual samples. Specifically, for each sample in the candidate datasets, we calculate a vulnerability score defined as:
$\text{Score}=\sum_{i=1}^{3}\mathbbm{1}^{(i,\text{before})}-\sum_{i=1}^{3}\mathbbm{1}^{(i,\text{after})},$ (1)
@@
where $\mathbbm{1}^{(i,\text{before})}$ and $\mathbbm{1}^{(i,\text{after}}$ indicate whether model $i$ (among the three evaluated: OpenAI o4-mini, Claude-3.7-Sonnet, Gemini-2.5-Flash) produce the correct answer before and after the gaslighting negation, respectively. Samples with higher scores indicate greater susceptibility to gaslighting and are prioritized during sampling. To ensure domain diversity, we stratify selection across categories defined by each dataset. As shown in Figure 2, for MMMU and CharXiv, we uniformly sample from non-overlapping subject categories. MathVista, however, presents a special case: its examples often span multiple overlapping skills (e.g., a single question tagged as both geometry and numeric commonsense). As a result, category counts in MathVista are not strictly balanced, though we aim for representative coverage across reasoning types. After careful curation to reflect both reasoning diversity and gaslighting vulnerability, GaslightingBench-R is concluded with 1,025 samples spanning 21 categories, comprising 400 from CharXiv, 300 from MMMU, and 325 from MathVista. Figure 3 illustrates several representative examples drawn from different categories within the dataset.
3.3 Performance Comparison
@@
MMMU [mmmu] MathVista [mathvista] CharXiv [wang2024charxiv]
@@
As shown in Table 1, we compare the performance of three state-of-the-art reasoning models, OpenAI-o4-mini, Claude-3.7-Sonnet and Gemini-2.5-Flash, before and after the introduction of gaslighting negation attacks across three multimodal reasoning benchmarks. All three models exhibit significant performance degradation after the introduction of gaslighting prompt. The average accuracy drop across the three datasets ranges from 25.6 to 28.8 percent. These results demonstrate that even the most advanced reasoning models struggle to maintain factual consistency and are unable to reliably defend correct answers under adversarial user interventions.
4 Result Analysis
Figure 4: Performance comparison of reasoning models on GaslightingBench-R before and after gaslighting negation attacks.
@@ 4 Result Analysis
Before gaslighting negation attack, OpenAI-o4-mini consistently outperforms its counterparts, achieving accuracies of 77.4 on MMMU, 77.1 on MathVista, and 65.2 on CharXiv, resulting in an overall average of 73.2%. Following the gaslighting negation attack, its performance drops to 52.1, 54.1, and 36.7, respectively, with an average of 47.6%. While this represents a substantial decline of 25.6%, it is the smallest average drop among the three evaluated models. In particular, Gemini-2.5-Flash achieves the highest pre-negation accuracy on MathVista (77.6), but suffers the largest performance decline (35.4) after gaslighting negation, highlighting its high vulnerability and limited robustness in maintaining correct mathematical reasoning under gaslighting negation attack.
@@
As shown in Figure 4, we further evaluate all three models on our newly constructed GaslightingBench-R. Compared to their performance on the existing benchmarks, the models exhibit substantially more severe degradation, with accuracy dropping by over 53% on average. In contrast, the average accuracy drops on MMMU, MathVista, and CharXiv range from 25.6% to 28.8%. This discrepancy highlights the unique difficulty of GaslightingBench-R, which is intentionally curated to target model susceptibility. Specifically, samples are selected based on their likelihood to induce belief reversals across multiple models, as described in Section 2, thereby maximizing their diagnostic value. Consequently, GaslightingBench-R serves as a focused and stress-tested benchmark for evaluating reasoning model robustness, offering deeper insights into belief collapse and behavioral inconsistency than conventional evaluations.
Figure 5: Accuracy comparison before and after gaslighting negation across subject categories in the MMMU benchmark. The bars represent model accuracy before (light) and after (dark) gaslighting negation, with the absolute accuracy drop annotated.
4.1 Category Analysis and Qualitative Result
@@
Category analysis. Figures 57 present a fine-grained breakdown of accuracy before and after gaslighting prompts for three reasoning models across thematic sub-categories in the MathVista, MMMU, and ChartXiv benchmarks. Figure 5 shows bar plots of model accuracy across six subject categories from the MMMU benchmark: Tech and Engineering, Humanities and Social Sciences, Health and Medicine, Science, Business, and Art and Design. Prior to gaslighting negation, all three models perform relatively well in Business, while exhibiting lower accuracy in Tech and Engineering. Following gaslighting negation attacks, the most substantial performance declines are observed (e.g., –37.3 for OpenAI-o4-mini, –30.0 for Gemini-2.5-flash), indicating that models are particularly fragile in domains requiring precise and high-stakes reasoning.
Figure 6: Reasoning skill-wise accuracy comparison on the MathVista benchmark before and after gaslighting negation attacks.
Figure 7: Subject-wise accuracy comparison on the CharXiv benchmark before and after gaslighting negation attacks.
In contrast, the Business category experiences the smallest declines on average (e.g., –16.6 for OpenAI-o4-mini, –18.0 for Claude-3.7-Sonnet), suggesting a relatively stable reasoning trajectory in economically contextual tasks. Figure 6 presents radar plots highlighting the three reasoning models’ performance across six core reasoning skills: geometry, arithmetic, algebraic, logical, statistical reasoning, and numeric commonsense. Before negation, all models exhibit strong performance in geometry and statistical reasoning. After negation, notable performance declines are observed in arithmetic reasoning and numeric commonsense, particularly for OpenAI-o4-mini and Gemini-2.5-Flash. suggesting these areas are particularly vulnerable to gaslighting prompts. These categories may be more susceptible to gaslighting due to their quantitative ambiguity and dependence on intermediate calculations [geva2020injecting]. Figure 7 provides a radar plot comparing model performance across eight scientific subjects in CharXiv: Mathematics, Physics, Quantitative Biology, Quantitative Finance, Statistics, Computer Science, Economics, and Electrical Engineering and Systems Science. Compared to MMMU and MathVista, while all three models suffer noticeable performance declines after gaslighting across every subject in CharXiv, the degradation is relatively uniform, showing the general vulnerability of gaslighting negation for the real-world scientific chart understanding.
Figure 8: Qualitative examples of gaslighting-induced answer reversals across reasoning models. The revised answers are accompanied by hallucinated rationales absent from the original explanation, revealing vulnerabilities in the reasoning process.
@@
Qualitative result. Figure 8 showcases qualitative examples of gaslighting-induced failures across the three reasoning models, highlighting the models’ behavior before and after exposure to a gaslighting negation attack. The top-left panel shows a math-based visual reasoning question, while the top-right involves a conceptual multimodal question about literary interpretation grounded in an image of a castle. In both cases, the models initially produce correct answers with coherent justifications, but a simple negation attack causes them to revise to incorrect responses. In the math example, all three models abandon the correct radius of 5 and switch to 4, despite the objective visual evidence. In the literary example, Claude-3.7-Sonnet and Gemini-2.5-Flash reverse their correct choice (“Castle Tales”) to the wrong option (“Retold Tale”), fabricating new rationales absent from their original reasoning. The bottom-left panel demonstrates a statistical reasoning question (“Is Cadet Blue the high median?”). Here, the models again answer correctly at first, but after negation, each shifts to an incorrect response while generating spurious explanations, revealing instability even in relatively straightforward yes/no reasoning tasks. The bottom-right panel highlights a measurement task requiring the use of a ruler. All three models initially compute the correct length (3 inches), yet under negation, they adjust their answers to incorrect values and rationalize the mistake with flawed geometric reasoning. Across all four examples, a consistent pattern emerges: reasoning models not only reverse correct answers when gaslighted but also hallucinate new justifications to support their revised positions. This behavior highlights a critical vulnerability, where models over-accommodate user feedback at the expense of internal consistency, demonstrating that explicit reasoning traces alone do not safeguard against belief manipulation.
5 Conclusion
@@
We have presented the first systematic investigation into the robustness of state-of-the-art reasoning models against adversarial gaslighting negation attacks. Despite their use of chain-of-thought reasoning and test-time scaling, models like OpenAI’s o4-mini, Claude-3.7-Sonnet and Gemini-2.5-Flash exhibit significant belief reversals when challenged with negation prompts after providing correct and well-justified answers. To diagnose this vulnerability more precisely, we introduce GaslightingBench-R, a curated benchmark targeting belief inconsistency in reasoning models. Our results show that GaslightingBench-R elicits even more noticeable failures than standard benchmarks, revealing a critical gap between reasoning transparency and belief stability. These findings call for rethinking evaluation protocols and advancing robustness strategies that go beyond correctness and interpretability to include resilience against adversarial conversational manipulations.
{credits}
5.0.1 Acknowledgements
This study was funded by the Science and Technology Commission of Shanghai Municipality (grant number 24511103100).